# Efficient optical beam displacement measurements using high-order Hermite-Gaussian modes

### v1

@@ Abstract
The lower bound for the uncertainty on the estimation of the parameter $g$, $\delta g$, is provided by the Cramér-Rao bound which establishes that $\delta g\geq 1/\sqrt{F(g)}$, where $F(g)$ is the Fisher information associated to the set of probabilities $P_{i}(g)$ of a given measurement, such that $F(g)=\sum_{i}1/P_{i}(g)\left[dP_{i}(g)/dg\right]^{2}$. The maximization of the Fisher information over all possible quantum measurements leads to the quantum Fisher information, and for a unitary operation, its value is given by $\mathcal{F}=4(\langle\hat{H}^{2}\rangle-\langle\hat{H}\rangle^{2})$, where $\hat{H}$ is the generator of the unitary transformation and the average is taken with respect to the initial quantum state [28]. Therefore, the quantum Fisher information for a coherent state with a mean photon number $N$ in the spatial mode of Eq.(1) is given by
$\mathcal{F}=N\langle u_{m}|\hat{x}^{2}|u_{m}\rangle=N(2m+1)w^{2}\,,$ (3)
from which we can evidence the metrological enhancement provided by the higher order modes once the uncertainty $\delta g$ is reduced by a $\sqrt{2m+1}$ factor.
By measuring the probabilities of detecting the photon state in the original mode $u_{m}$ ($P_{m}$) or in the populated modes $u_{m+1}$ and $u_{m-1}$ ($P^{\prime}_{m}$), we get the following probabilities from Eq.(2): $P_{m}=1-(2m+1)g^{2}w^{2}/4$ and $P^{\prime}_{m}=(2m+1)g^{2}w^{2}/4$. The Fisher information arising from this (binary) measurement with $N$ uncorrelated photons from a coherent state is given by
@@
Hence, the proposed measurement is efficient under the small tilt approximation.
For the lateral displacement $d$, the unitary operator is given by $\exp{(-id\hat{p})}$, where
$\hat{p}=-i\frac{\partial}{\partial x}\,.$ (5)
Thus, the quantum Fisher information for this process is given by
$\mathcal{F}=N\langle u_{m}|\hat{p}^{2}|u_{m}\rangle=\frac{4N(2m+1)}{w^{2}}\,.$ (6)
@@
$e^{-id\hat{p}}u_{m}\approx u_{m}+\frac{d}{w}\left(\sqrt{m+1}\,u_{m+1}-\sqrt{m}\,u_{m-1}\right)\,,$ (7)
@@
and the probabilities associated with measuring the photon in the original state ($u_{m}$) and the populated ones ($u_{m+1}$ and $u_{m-1}$) are given by $P_{m}=1-(2m+1)d^{2}/w^{2}$ and $P^{\prime}_{m}=(2m+1)d^{2}/w^{2}$. The corresponding Fisher information with $N$ uncorrelated photons is then
$F(g)=\frac{4N(2m+1)}{w^{2}}+\mathcal{O}(d^{2}/w^{2})\,,$ (8)
@@
showing once again the efficiency in the proposed measurement protocol.
In order to realize such measurements, we used a Mach-Zehnder interferometer with an additional mirror in one arm. When the longitudinal phase difference in the two arms are equivalent to zero ($\varphi=0$), the odd spatial modes acquires a $\pi$ phase difference with respect to the even spatial modes due to the extra reflection in one arm, such that the bright port for the odd modes corresponds to the dark port for the even modes [29]. Therefore, the modes are separated by their parity, and hence, the mode intensities regarding the second and third terms in the expansions (2) and (7), which are proportional to the probabilities $P_{m}$ and $P^{\prime}_{m}$, are obtained by monitoring the output ports of the interferometer in the situation where $\varphi=0$. In fact, due to the conservation of energy, knowing the total power of the input beam, it is sufficient to monitor only one port of the interferometer, to measure the increase (decrease) in optical power in the dark (bright) port.
$I=\frac{I_{tot}}{2}\left(1+\nu_{m}\cos\varphi\right)\,,$ (9)
where $I_{tot}$ is the total input optical power, and $\nu_{m}=|(|E_{o}|^{2}-|E_{e}|^{2})|/(|E_{o}|^{2}+|E_{e}|^{2})=(|I_{o}-I_{e}|)/(I_{o}+I_{e})$ is the visibility of the interferometer for the mode $m$, which reveals the contrast between the different parity modes. For a well-defined input parity mode ($u_{m}$) the visibility is equal to 1. However, for displaced beams with respect to the optical axis of the interferometer, the visibility will be different from 1, once they can be seen as the expressions (2) and (7), which present a small component in the opposite parity. Since the optical power in a given mode is proportional to the probabilities of individual photons, we can write that $\nu_{m}=1-2P^{\prime}_{m}$, such that
$\nu_{m}^{(g)}$ $=$ 1($2\displaystyle 1-(2m+1)g^{2}w^{2}/$ 2 (10)
$\nu_{m}^{(d)}$ $=$ $1-2(2m+1)d^{2}/w^{2}\,,$ (11)
for the angular and lateral displacement, respectively.
Applying an external modulation $\Omega$ in the displacement parameters such that $g\to g_{0}+g\sin(\Omega t)$ and $d\to d_{0}+d\sin(\Omega t)$, where $g_{0}$ and $d_{0}$ account for an unavoidable experimental offset, and substituting Eqs.(10) and (11) into (9) with their respective modulations, we end up with
$I^{(d)}=\frac{I_{tot}}{2}\Big{[}\Gamma_{0}-4(2m+1)\frac{d_{0}d}{w^{2}}\cos\varphi\sin(\Omega t)$
$+(2m+1)\frac{d^{2}}{w^{2}}\cos\varphi\cos(2\Omega t)\Big{]}\,,$ (12)
$I^{(g)}=\frac{I_{tot}}{2}\Big{[}\Gamma^{\prime}_{0}-(2m+1)g_{0}gw^{2}\cos\varphi\sin(\Omega t)$
$+(2m+1)\frac{g^{2}w^{2}}{4}\cos\varphi\cos(2\Omega t)\Big{]}.$ (13)
Therefore, we see that there are two spectral components ($\Omega$ and $2\Omega$) that carry information about the displacement, which will be maximized at $\varphi=\{0,\pi\}$, precisely when the interferometer acts as a parity selector. The frequency $2\Omega$ will be most relevant, since its magnitude depends only on the amplitude of the modulation, unlike the frequency component $\Omega$. Moreover, they exhibit the magnification factor of $(2m+1)$ that characterizes the metrological enhancement present in Eqs.(3) and (6).
Experiment.—The experimental setup is depicted in Fig.(1). A spatial light modulator (Hamamatsu LCOS-SLM X10468) is used to generate the spatial modes $u_{m}$  [^1] We utilized the repository [31] to generate the necessary holograms for producing the transverse modes from a $633$ nm He-Ne laser beam. The generated beam is collimated by the lenses $L_{1}$ and $L_{2}$ to a width of $w=590\,\mu m$. Two piezoelectric actuators are attached to the mirror mount just before the interferometer: the $PZT_{L}$ is attached to the back of the mirror to ensure a pure lateral displacement of the beam; the $PZT_{A}$ is screw actuator (Thorlabs Polaris-P20A) installed on a mirror mounting barrel to ensure an angular displacement. On one of the interferometer mirrors, we coupled a third piezoelectric actuator ($PZT$) which allows us to fine-tune the relative phase $\varphi$ of the interferometer. The output ports of the interferometer are monitored by a CCD camera, to visually check the interference quality, and a photo-detector (Thorlabs PDA100A2).
Figure 1: Experimental setup. A spatial light modulator (SLM) prepares the desired $u_{m}$ mode. The signal from the photo-detector (PD) is demodulated in a lock-in amplifier (Liquid Instruments, Moku:Pro).
Initially, we align the interferometer with the incoming laser beam in order to ensure the highest possible visibility, and we consider this as our starting point. The visibility we obtained for the Gaussian ($u_{0}$) mode is $97\%$. Since high-order modes are more sensitive to misalignment, we eventually have to fine-tune the alignment in one of the interferometer mirrors when dealing with these modes. After the initial alignment procedures, we apply the displacement (lateral or angular) modulation with peak-to-peak drive voltage of $0.8$ V at a frequency $\Omega=650$ Hz, well bellow the mechanical resonance of the piezos. The measured signal is demodulated in a lock-in amplifier at frequency $2\Omega$ in order to measure the amplitude of the applied modulation. As seen from Eqs.(12) and (13) the demodulated signal is proportional to $\cos\varphi$, which implies that it is maximized at the constructive (destructive) output port, when $\varphi=0$ ($\varphi=\pi$). Therefore, we experimentally drive the interferometer piezoactuator (PZT) in order to maximize the demodulated signal for each incoming transverse mode $u_{m}$. Ideally, the interferometer would be phase-locked at this position, but we have verified that the phase is quite stable during the measurements for each transverse mode.
The experimental results for the angular displacement is depicted in Fig.(2). It shows the amplitude of the demodulated signal at frequency $2\Omega$ as a function of the input spatial mode for a given modulation amplitude $g$. The gain in the lock-in amplifier was set to $80$ dB. In order to capture the dependency solely with the mode order $m$, we fixed the optical power to $P=2\,\mu W$, except for modes $m=\{16,17,18,19,20\}$, which were taken at $P=1.5\,\mu W$, in order to avoid $1\,V$ saturation signal at the lock-in amplifier. However, we carefully rescaled their values owing to its proportionality to the input optical power.
We note from Eq.(13), that the amplitude at frequency $2\Omega$ is linearly proportional to the mode order $m$, as $V_{2\Omega}=V_{0}(2m+1)$, where $V_{0}$ is the signal value for the Gaussian mode. Thus, to check the dataset trend, we performed a linear fit with $V_{0}$ as the free parameter. The obtained value $32.35$ mV agrees with the experimental value $(33\pm 2)$ mV , where the error was taken from the random fluctuations over repeated measurements.
This result shows the magnification coming from the Hermite-Gauss mode order $m$, which amplifies the signal but not the noise. The overall effect is the increase in the signal-to-noise (SNR), which allows for detection of smaller tilt values. Regarding Eq.(13), we see the estimated value of $g$ is inversely proportional to $\sqrt{2m+1}$, which attest the metrological gain deriving from the quantum Fisher information (3). In fact, our setup was able to achieve a magnification factor of $41$ ($m=20$) in the SNR, which accounts for a improvement factor of $6.4$ in the estimability of the tilt. This result would correspond to a $16$ dB squeezed light approach in the $HG_{1,0}$ mode, a value not reached so far, to the best of our knowledge.
@@
Figure 2: Figure (a) presents the results for angular displacement, while figure (b) shows the results for lateral displacement. In both figures, blue dots represent the experimental measurements, and the red lines indicate linear fits to our data. The error bar is small and, therefore, it isn’t represented in the figure.
The results concerning the lateral displacement are shown in Fig.(2). The same analysis was realized for the data set as in the angular displacement. The input power was slightly lower for modes $m=\{15,16,17,18,19,20\}$, due to the low SLM efficiency for such high-order modes, but proper rescale was taken, as in the previous case. The fitting result for the $V_{0}$ parameter was $22.25$ mV, which matches with the experimental value $(23\pm 2)$ mV, showing good agreement with the theory.
Conclusion.—We have described an interferometric technique to measure lateral and angular displacements of an optical beam, which has relevant technological applications, as atomic force microscopy, for instance. The technique relies on the enhanced metrological performance provided by the Hermite-Gaussian modes. We were able to improve the SNR by a factor of $41$ compared to the Gaussian mode in both displacements. One major limitation is the production of such high-order modes, which tends to lose efficiency as $m$ increases; moreover, since the mode width increases with the order, bigger optics are required, which imposes limitations on our small setup. On the other hand, there is no need for a mode-matched local oscillator with a higher intensity or a multimode squeezed state of light, which, to our point of view, makes the technique more practical.
Acknowledgements.—The authors would like to acknowledge the financial support from the Brazilian agencies: Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES), Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) and Instituto Nacional de Ciência e Tecnologia de Informação Quântica (INCT-IQ 465469/2014-0). A. L. S. S. J. acknowledges financial support from Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP).

### v2

@@ A Hermite-Gaussian mode can then be written as
By measuring the probabilities of detecting the photon state in the original mode $HG_{m,0}$ ($P$) or in the populated modes $HG_{m+1,0}$ and $HG_{m-1,0}$ ($P^{\prime}$), we get the following probabilities from Eq.(Optimal beam displacement measurements using high-order structured light modes): $P=1-(2m+1)g^{2}w^{2}/4$ and $P^{\prime}=(2m+1)g^{2}w^{2}/4$. The Fisher information arising from this (binary) measurement with $N$ uncorrelated photons from a coherent state is given by
@@
Hence, the proposed measurement is optimal under the small tilt approximation.
Consider now the case of a small tilt in a Laguerre-Gauss beam. The quantum Fisher information for this process is given by
$\mathcal{F}=N\langle LG_{m,n}|\hat{x}^{2}|LG_{m,n}\rangle=N(m+n+1)w^{2}\,.$ (31)
@@
$e^{-ig\hat{x}}\ket{LG_{m,n}}\approx(1-ig\hat{x})\ket{LG_{m,n}}=\ket{LG_{m,n}}$
$-\frac{igw}{2\sqrt{2}}\Big{(}\sqrt{m+1}\,\ket{LG_{m+1,n}}+\sqrt{m}\,\ket{LG_{m-1,n}}$
$+\sqrt{n+1}\,\ket{LG_{m,n+1}}+\sqrt{n}\,\ket{LG_{m,n-1}}\Big{)}\,,$ (32)
@@
and the probabilities associated with measuring the photon in the original state ($LG_{m,n}$) and the populated ones ($LG_{m+1,n}$, $LG_{m-1,n}$, $LG_{m,n+1}$ and $LG_{m,n-1}$) are given by $P=1-(m+n+1)g^{2}w^{2}/4$ and $P^{\prime}=(m+n+1)g^{2}w^{2}/4$. The corresponding Fisher information with $N$ uncorrelated photons is then
$F(g)=N(m+n+1)w^{2}+\mathcal{O}(g^{2}w^{2})\,,$ (33)
@@
showing once again that the proposed measurement protocol is optimal.
Finally, the case of a lateral displacement of a Laguerre-Gauss beam is accounted by the quantum Fisher information:
$\mathcal{F}=4N\langle LG_{m,n}|\hat{p}_{x}^{2}|LG_{m,n}\rangle=\frac{4N(m+n+1)}{w^{2}}\,.$ (34)
The first-order expansion of the displaced Laguerre-Gauss mode is given by:
$e^{-id\hat{p}_{x}}\ket{LG_{m,n}}\approx(1-id\hat{p}_{x})\ket{LG_{m,n}}=\ket{LG_{m,n}}$
$+\frac{d}{\sqrt{2}w}\left(\sqrt{m+1}\ket{LG_{m+1,n}}-\sqrt{m}\ket{LG_{m-1,n}}\right.$
$\left.+\sqrt{n+1}\ket{LG_{m,n+1}}-\sqrt{n}\ket{LG_{m,n-1}}\right)\,,$ (35)
such that the probabilities associated with measuring the original state ($LG_{m,n}$) and the populated ones ($LG_{m+1,n}$, $LG_{m,n+1}$, $LG_{m-1,n}$ and $LG_{m,n-1}$) are given by $P=1-(m+n+1)d^{2}/w^{2}$ and $P^{\prime}=(m+n+1)d^{2}/w^{2}$. The corresponding Fisher information with $N$ uncorrelated photons is then
$F(d)=\frac{4N(m+n+1)}{w^{2}}+\mathcal{O}(d^{2}/w^{2})\,,$ (36)
showing once again that the proposed measurement protocol is optimal.
The results the angular displacement for the Hermite-Gauss modes are shown in the (S1) (a). In order to capture the dependence on the mode order $m$, we fixed the optical power at $P=2.0\,\mu W$, except for modes with order higher than $m=10$, which were measured with $P=1.1\,\mu W$ to avoid the $1\,V$ saturation in the lock-in amplifier. We carefully rescaled these values owing to its proportionality with the input optical power. This procedure was applied to all results, and we have provided detailed tables below with specific values for each set of results. The fitting result for the $V_{0}$ parameter was $50$ mV, which matches with the experimental value (48±2(48$\pm$ 2( 48 ± 2) mV, showing good agreement with the theory.
In Figure S1 (b), we illustrate the case where $l=0$, leading to $m=n=p$ for lateral displacement. For this configuration, the linear fit is performed with $V_{2\Omega}=V_{0}(2p+1)$, resulting in a fitted value of $25$ mV, compared to the experimental result of ($28\pm 2$) mV.
In the tables below, we present the experimental results, with the error bars, obtained directly from the Lock-in amplifier. All modes were carefully calibrated to have the same power $P$ ($\mu$W), as indicated in the tables, ensuring data consistency. Due to the saturation of the Lock-in amplifier at 1.1 V, it was necessary to recalibrate to a lower power for the HGm,0 modes in angular displacement with $m$ greater than 10. Therefore, we rescaled the obtained values, considering they are proportional to the input power.
@@
Figure S1: Figure (a) presents the results for angular displacement using Hermite-Gaussian modes $HG_{m,0}$, while figure (b) shows the results for Laguerre-Gaussian modes with $l=0$ and $m=n=p$, for lateral displacement. In both figures, the blue dots represent the experimental measurements, whereas the red lines correspond to linear fits to the obtained data. The error bars are small and, therefore, are not displayed in the figures.
Table S1: Using Hermite-Gaussian modes
$m$ $V_{2\Omega}$ (mW) P ($\mu$W)
0 $25\pm 2$
1 $83\pm 2$
2 $124\pm 2$
3 $165\pm 2$
4 $234\pm 2$
5 $265\pm 2$
6 $299\pm 2$
7 $367\pm 2$
8 $407\pm 4$
9 $464\pm 4$
10 $507\pm 4$ $2.0\pm 0.1$
11 $566\pm 4$
12 $609\pm 4$
13 $676\pm 4$
14 $706\pm 4$
15 $758\pm 4$
16 $824\pm 7$
17 $869\pm 8$
18 $965\pm 8$
19 $932\pm 8$
20 $1038\pm 8$
Table S1: Lateral displacement
$m$ $V_{2\Omega}$ (mW) P ($\mu$W)
0 $48\pm 4$
1 $136\pm 4$
2 $249\pm 4$
3 $353\pm 4$
4 $410\pm 4$
5 $525\pm 4$ $2.0\pm 0.1$
6 $647\pm 6$
7 $799\pm 8$
8 $860\pm 8$
9 $960\pm 10$
10 $1062\pm 10$
11 $593\pm 4$
12 $619\pm 4$
13 $688\pm 4$
14 $772\pm 6$
15 $796\pm 6$ $1.0\pm 0.1$
16 $815\pm 8$
17 $861\pm 8$
18 $918\pm 8$
19 $946\pm 10$
20 $1066\pm 10$
Table S2: Angular displacement
Table S3: Using Laguerre-Gaussian modes
$p$ $V_{2\Omega}$ (mW) P ($\mu$W)
0 $29\pm 4$
1 $83\pm 4$
2 $126\pm 4$
3 $179\pm 4$
4 $221\pm 6$
5 $283\pm 6$ $1.0\pm 0.1$
6 $331\pm 6$
7 $412\pm 6$
8 $449\pm 8$
9 $473\pm 8$
10 $501\pm 8$
Table S3: With $\ell=0$
$\ell$ $V_{2\Omega}$ (mW) P ($\mu$W)
0 $26\pm 4$
1 $54\pm 4$
2 $77\pm 4$
3 $99\pm 4$
4 $127\pm 4$
5 $159\pm 4$
6 $170\pm 6$
7 $201\pm 6$
8 $228\pm 6$
9 $244\pm 6$
10 $278\pm 6$ $1.0\pm 0.1$
11 $284\pm 8$
12 $309\pm 8$
13 $328\pm 8$
14 $345\pm 8$
15 $370\pm 8$
16 $406\pm 10$
17 $430\pm 10$
18 $477\pm 10$
19 $487\pm 10$
20 $516\pm 10$
Table S4: With $p=0$